OpenStudy (shelby1290):

Subtract the sum of 3x^2-2x+7 and 4x^2-x-12 FROM the sum of 5x+2x^2-6 and 3x^2-5

Vocaloid (vocaloid):

(5x+2x^2-6 + 3x^2-5) - (3x^2-2x+7 + 4x^2-x-12) it looks messy right now, but start by combining like terms within the parentheses

OpenStudy (shelby1290):

Wait, why did you switch the order? Isn't it supposed to be (3x^2-2x+7 + 4x^2-x-12) - (5x+2x^2-6 + 3x^2-5)

Vocaloid (vocaloid):

"subtracting A FROM B" means B - A

OpenStudy (shelby1290):

isn't it A - B ?

Vocaloid (vocaloid):

I know it's confusing, but that's the what the word "from" means when we're subtracting

OpenStudy (shelby1290):

okay

Vocaloid (vocaloid):

"I want to take away 3 from 5" means 5 - 3, agreed?

OpenStudy (shelby1290):

ohh okay i see

Vocaloid (vocaloid):

good, so start by combining like terms

Vocaloid (vocaloid):

just focus on the first parentheses part (3x^2-2x+7 + 4x^2-x-12) we can combine the x^2 terms and the x terms together

OpenStudy (shelby1290):

5x+2x^2-6+3x^2-5 =2x^2+3x^2+5x-6-5 =5x^2+5x-11

Vocaloid (vocaloid):

good, now leave that alone and work on the second part

OpenStudy (shelby1290):

3x^2-2x+7+4x^2-x-12 =3x^2+4x^2-2x-x+7-12 =7x^2-3x-5

Vocaloid (vocaloid):

good. that gives us (5x^2+5x-11) - (7x^2-3x-5) now distribute the negative sign in the middle of the two parentheses

OpenStudy (shelby1290):

(5x^2+5x-11) - (7x^2-3x-5) =5x^2+5x-11-7x^2-3x-5 =5x^2-7x^2+5x-3x-11-5 =-2x^2+2x-16

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

because of the negative sign, each sign in the second parentheses switches

Vocaloid (vocaloid):

(5x^2+5x-11) - (7x^2-3x-5) = (5x^2+5x-11) - 7x^2 + 3x + 5

Vocaloid (vocaloid):

(this is part of the distributive property)

OpenStudy (shelby1290):

alright so... (5x^2+5x-11) - 7x^2 + 3x + 5 =5x^2-7x^2+5x+3x-11+5 =-2x^2+8x-6

Vocaloid (vocaloid):

good

OpenStudy (shelby1290):

thank you